Find a power series solution pf xy'=y

1
Expert's answer
2022-02-08T17:56:01-0500

CORRECTED SOLUTION.

Let "y=\sum\limits_{n=0}^{+\infty}c_nx^n". Then "xy'=x\sum\limits_{n=0}^{+\infty}nc_nx^{n-1}=\sum\limits_{n=0}^{+\infty}nc_nx^n".

Since "xy'=y", it must be "nc_n=c_n" for all "n=0,1,2,\dots"

This means "c_n=0" for all "n\ne 1" and "c_1" can be arbitrary.


Answer. "y=c_1x".